Can you help me...
A) Find the average rate of change of the function

f(x)= 1/x+5; x=5, x=15
b) Find the maximum or minimum value of the function
f(x) = 1+5x+5x^2

Answer by stanbon(48535)

A) Find the average rate of change of the function
f(x)= 1/x+5; x=5, x=15
ave. rate of change = [f(15)-f(5)]/[15-5]
=[1/20 - 1/10]/[10]
=[-1/20]/10
=[-1/200]
---------------------
b) Find the maximum or minimum value of the function
f(x) = 1+5x+5x^2
min occurs at x=-b/2a= -5/10=-1/2
f(-1/2)=1-(5/2)+(5/4)
=-1/4
min occurs at (-1/2,-1/4)
Cheers,
Stan H.